OpenStudy (anonymous):

factor and use identities to simplify: sin(t)xcos^2(t)-sin(t)

OpenStudy (anonymous):

this might be easier to read \[\sin \left( t \right)\times \cos ^{2}\left( t \right)-\sin (t)\]

OpenStudy (turingtest):

first note that\[\sin^2t+\cos^2t=1\rightarrow \cos^2t-1=-\sin^2t\]we now see that\[\sin(t)\cos^2t-\sin(t)=\sin t(\cos^2t-1)=\sin t(-\sin^2t)=-\sin^3t\]

OpenStudy (anonymous):

so you just factored out the sin(t) and cos^2(t)-1 = -sin^2(t)?

OpenStudy (turingtest):

yup

OpenStudy (anonymous):

ok thanks